OpenStudy (anonymous):

Find the value of x in the equation 2√x + 20 = 15 + 3√x .

OpenStudy (anonymous):

\[2\sqrt{x}+20=15+3\sqrt{x}\]

OpenStudy (anonymous):

getting sqrt(x) on one side and constant on the other gives sqrt(x) = 5. Can u understand that?

OpenStudy (anonymous):

Then square both sides to get x = 25.

OpenStudy (anonymous):

\[\sqrt{x}+5=\sqrt{x+45} \] what's x? ok another example

OpenStudy (ash2326):

square both sides \[{(\sqrt x+5)}^2=\sqrt{x+45}^2\] we'll get \[ x+ 25+10 \sqrt x= x+45\] cancel both x we get \[10\sqrt x=20\] \[\sqrt x=2\] \[x=4\]

OpenStudy (anonymous):

thank you for showing all steps